# footnotetext: $\dagger$ First authors

### v1

@@ Abstract
Biological and artificial neural networks develop internal representations that enable them to perform complex tasks. In artificial networks, the effectiveness of these models relies on their ability to build task specific representation, a process influenced by interactions among datasets, architectures, initialization strategies, and optimization algorithms. Prior studies highlight that different initializations can place networks in either a lazy regime, where representations remain static, or a rich/feature learning regime, where representations evolve dynamically. Here, we examine how initialization influences learning dynamics in deep linear neural networks, deriving exact solutions for $\lambda$-balanced’ initializations—defined by the relative scale of weights across layers. These solutions capture the evolution of representations and the Neural Tangent Kernel across the spectrum from the rich to the lazy regimes. Our findings deepen the theoretical understanding of the impact of weight initialization on learning regimes, with implications for continual learning, reversal learning, and transfer learning, relevant to both neuroscience and practical applications.
@@ 1 Introduction
Biological and artificial neural networks learn internal representations that enable complex tasks such as categorization, reasoning, and decision-making. Both systems often develop similar representations from comparable stimuli, suggesting shared information processing mechanisms Yamins et al. (2014). This similarity, though not fully understood, has drawn interest from neuroscience, AI, and cognitive science Haxby et al. (2001); Laakso & Cottrell (2000); Morcos et al. (2018); Kornblith et al. (2019); Moschella et al. (2022). The success of neural models relies on their ability to form these representations and extract relevant features from data to build internal representations, a complex process that in machine learning is defined by two regimes: lazy and rich Saxe et al. (2014); Pennington et al. (2017); Chizat et al. (2019); Bahri et al. (2020).
@@
Lazy regime. Extensive research has identified a fundamental phenomenon in overparameterized neural networks: during training, these networks frequently remain near their linearized form, undergoing minimal changes in the parameter space Chizat et al. (2019). Consequently, they adopt learning dynamics akin to kernel regression, characterized by the Neural Tangent Kernel (NTK) matrix and exhibiting exponential learning behavior Du et al. (2018); Jacot et al. (2018); Du et al. (2019); Allen-Zhu et al. (2019a, b); Zou et al. (2020). This behavior, known as the lazy or kernel regime, typically occurs in infinitely wide architectures and can be triggered by large variance initialization at the start of training Jacot et al. (2018); Chizat et al. (2019). While the lazy regime offers valuable insights into how networks converge to a global minimum, it does not fully account for the generalization capabilities of neural networks trained with standard initializations. It is, therefore, widely believed that another regime, driven by small or vanishing initializations, underpins some of the successes of neural networks.
@@
Rich regime. In contrast, the rich feature-learning regime is characterized by a NTK that evolves throughout training, accompanied by non-convex dynamics that navigate saddle points Baldi & Hornik (1989); Saxe et al. (2014, 2019b); Jacot et al. (2021). This regime features sigmoidal learning curves and simplicity biases, such as low-rankness Li et al. (2020) or sparsity Woodworth et al. (2020). Numerous studies have shown that the absolute scale of initialization drives the rich regime, which typically emerges at small initialization scales Chizat et al. (2019); Geiger et al. (2020). However, it’s also been shown that even at small initialization scales, differences in weight magnitudes between layers can induce the lazy learning regime Azulay et al. (2021); Kunin et al. (2024). This highlights the significance of both absolute scale (initialization variance) and relative scale (difference in weight magnitude between layers) in generating diverse learning dynamics. Beyond absolute scale and relative scale, additional aspects of initialization can profoundly affect feature learning, including the effective rank of the weight matrices Liu et al. (2023), layer-specific initialization variances Yang & Hu (2020); Luo et al. (2021); Yang et al. (2022), and the use of large learning rates Lewkowycz et al. (2020); Ba et al. (2022); Zhu et al. (2023); Cui et al. (2024). These findings illustrate the effect of initialization on inducing complex learning behavior through the resulting dynamics. Here we develop a solvable model which captures these diverse phenomena.
Rich and lazy regimes in the brain. The distinction between rich and lazy learning may also hold implications for neuroscience, where neural representations have been argued to have task-specific or task-agnostic characteristics in different settings Farrell et al. (2023a); Ostojic & Fusi (2024); Tye et al. (2024). The lazy regime can be linked to the non-linear mixed selectivity of neurons, where task variables are represented in a high-dimensional space which mixes various potentially relevant variables Raposo et al. (2014); Tang et al. (2019); Rigotti et al. (2013); Bernardi et al. (2020). Conversely, the rich regime aligns with linear mixed selectivity Tye et al. (2024) and the manifold learning regime, where the brain encodes tasks on a structured, low-dimensional, task-specific manifold, as observed in grid cells within the entorhinal cortex Chaudhuri et al. (2019); Bernardi et al. (2020); Flesch et al. (2022).
@@
Figure 1: A minimal model of the rich and lazy regimes. A We examine a deep and wide linear network trained using gradient descent starting from an initialization characterized by a relative scale parameter $\lambda$— which characterizes the difference in the weight covariance between the first and second layers. B Network output for an example task over training time, starting from a range of relative scale values. The dynamics are influenced by the initialization. Solid lines represent simulations, while dotted lines indicate the analytical solutions derived in this work. C LeCun weight initialization LeCun et al. (1998) in the limit matches a $\lambda$ balanced weights. Plot of ${\bf W}_{2}(0)^{T}{\bf W}_{2}(0)-{\bf W}_{1}(0){\bf W}_{1}(0)^{T}$ for a two layer network initialized with LeCun initialization
@@
Despite significant advances, these learning regimes and their characterization are not yet fully understood and would benefit from clearer theoretical predictions, particularly regarding the influence of prior knowledge (initialization) on the learning regime. In this work, we address this gap by deriving exact solutions for the learning dynamics in deep linear networks as a function of network initialization, providing one of the few analytical models of the rich and lazy regimes in wide and deep neural networks Xu & Ziyin (2024); Kunin et al. (2024); Tu et al. (2024). To illustrate the dramatic effect of initialization and the kind of phenomenon we build a theory for, we consider a two-layer linear network parameterized by an encoding layer ${\bf W}_{1}$ and a decoding layer ${\bf W}_{2}$ (Fig.1A). This network can be initialized with different relative scalings, such that ${\bf W}_{1}{\bf W}_{1}^{T}\succ{\bf W}_{2}^{T}{\bf W}_{2}$, ${\bf W}_{1}{\bf W}_{1}^{T}\ \prec{\bf W}_{2}^{T}{\bf W}_{2}$, or ${\bf W}_{1}{\bf W}_{1}^{T}={\bf W}_{2}^{T}{\bf W}_{2}$, while maintaining the same absolute scale. As shown in Fig.1B, the choice of relative scaling can result in drastically different learning trajectories and representations and the theory we develop over the course of this paper describes these effects. Through these solutions, we aim to gain insights into the rich and lazy regimes, as well as the transition between them during training, by examining the impact of relative scaling. As shown in Fig.1C, initialization methods used in practice, such as LeCun initialization in wide networks, approximate the relative scaling initialization explored in this paper, making it relevant to machine learning community as further demonstrated in Kunin et al. (2024). We consider applications relevant to machine learning and neuroscience, including continual learning Kirkpatrick et al. (2017); Zenke et al. (2017); Parisi et al. (2019), reversal learning Erdeniz & Atalay (2010) and transfer learning Taylor & Stone (2009); Thrun & Pratt (2012); Lampinen & Ganguli (2018); Gerace et al. (2022).
@@
  • We derive explicit solutions for the gradient flow, internal representational similarity, and finite-width NTK in unequal-input-output two-layer deep linear networks, under a broad range of lambda-balanced initialization conditions (Section 4).
@@
  • We present applications of these solutions relevant to both the neuroscience and machine learning field, providing exact solutions for continual learning dynamics, reversal learning dynamics, and transfer learning (Section 6).
@@ 2 Related Work
Linear networks. Our work builds upon a rich body of research on deep linear networks, which, despite their simplicity, have proven to be valuable models for understanding more complex neural networks Baldi & Hornik (1989); Fukumizu (1998); Saxe et al. (2014). Previous research has extensively analyzed convergence Arora et al. (2018a); Du & Hu (2019), generalization properties Lampinen & Ganguli (2018); Poggio et al. (2018); Huh (2020), and the implicit bias of gradient descent Arora et al. (2019a); Woodworth et al. (2020); Chizat & Bach (2020); Kunin et al. (2022) in linear networks. These studies have also revealed that deep linear networks have intricate fixed point structures and nonlinear learning dynamics in parameter and function space, reminiscent of phenomena observed in nonlinear networks Arora et al. (2018b); Lampinen & Ganguli (2018). Seminal work by Saxe et al. (2014) laid the groundwork by providing exact solutions to gradient flow dynamics under task-aligned initializations, demonstrating that the largest singular values are learned first during training. This analysis has been extended to deep linear networks Arora et al. (2018b, 2019a); Ziyin et al. (2022) with more flexible initialization schemes Gidel et al. (2019); Tarmoun et al. (2021); Gissin et al. (2019). This work directly builds on the matrix Riccati formulation proposed by Fukumizu (1998); Braun et al. (2022) which extends these solutions to wide networks. We extend and refine these results to obtain the dynamics for lambda-balanced initialization dynamics of networks to more clearly demonstrate the impact of initialization on rich and lazy learning regimes also developed in Tu et al. (2024) for a set of orthogonal initalizations. Our work extends previous analysis Xu & Ziyin (2024); Kunin et al. (2024) of these regime to wide networks. Previous studies leveraged these solutions primarily to characterize convergence rates; however, our work goes beyond this by providing a comprehensive characterization of the complete dynamics of the system Tarmoun et al. (2021).
@@
Infinite-width networks. Recent advances in understanding the rich regime have largely stemmed from examining how the initialization variance and layer-wise learning rates must scale in the infinite-width limit to maintain consistent behavior in activations, gradients, and outputs. Several studies have employed statistical mechanics tools to derive analytical solutions for the rich population dynamics of two-layer nonlinear neural networks initialized using the mean field parameterization Mei et al. (2018); Rotskoff & Vanden-Eijnden (2018); Chizat & Bach (2018); Sirignano & Spiliopoulos (2020); Rotskoff & Vanden-Eijnden (2022); Sirignano & Spiliopoulos (2020). Other methods for analyzing deep network dynamics include the NTK limit, where the network effectively performs kernel regression without feature learning Jacot et al. (2018); Lee et al. (2019); Arora et al. (2019b). Our solution allows us to the study the evolution of the NTK and the influence of absolute scale and relative scale on the transition between lazy and rich learning in finite width networks Jacot et al. (2021); Xu & Ziyin (2024); Kunin et al. (2024); Chizat et al. (2019). Furthermore, these approaches typically require numerical integration or operate within a limited learning regime, and are unable to describe the learning dynamics of hidden representations. Instead, our work focuses on the impact of initialization on representation learning dynamics and derives explicit analytical solutions within tractable models.
@@ 3 Preliminaries
and the initialization ${\bf W}_{2}(0)$,${\bf W}_{1}(0)$. Our objective is to describe the entire dynamics of the network’s output and internal representations based on this initialization and the task statistics. We consider an approach first introduced in the foundational work of Fukumizu Fukumizu (1998) and extended in recent work by Braun et al. (2022), which rather than consider the dynamics of the parameters directly, we consider the dynamics of a matrix of the important statistics. In particular, defining ${\bf Q}=\begin{bmatrix}{\bf W}_{1}&{\bf W}_{2}^{T}\end{bmatrix}^{T}\in\mathbb{R}^{(N_{i}+N_{o})\times N_{h}}$, we consider the $(N_{i}+N_{o})\times(N_{i}+N_{o})$ matrix
@@
where $+$ denotes the pseudoinverse; and the network’s finite-width NTK Jacot et al. (2018); Lee et al. (2019); Arora et al. (2019b)
@@
where ${\bf I}$ is the identity matrix and $\otimes$ is the Kronecker product. Hence, the dynamics of ${\bf Q}{\bf Q}^{T}$ describes the important aspects of network behaviour.
@@ 4 Exact Learning Dynamics
In this section, we derive an exact solution for ${\bf Q}{\bf Q}^{T}$ providing a clean understanding of the learning dynamics, convergence behavior, and generalization properties of two-layer linear networks with prior knowledge.
@@
  • A1 (Whitened input). The input data is whitened, that is ${\bf\tilde{\Sigma}}^{xx}={\bf I}$.
  • A2 (Lambda-balanced). The network’s weight matrices are lambda-balanced at the beginning of training, that is ${\bf W}_{2}(0)^{T}{\bf W}_{2}(0)-{\bf W}_{1}(0){\bf W}_{1}(0)^{T}=\lambda\mathbf{I}$. If this condition holds at initialization, it will persist throughout training Saxe et al. (2014); Arora et al. (2018a). For completeness, we prove this in Appendix A.
@@
  • A4 (Full-rank). The input-output correlation of the task and the initial state of the network function have full rank, that is $\operatorname{rank}({\bf\tilde{\Sigma}}^{xy})=\operatorname{rank}({\bf W}_{2}(0){\bf W}_{1}(0))=\min(N_{i},N_{o})$.
These assumptions are strictly weaker than prior works by Fukumizu (1998); Braun et al. (2022); Kunin et al. (2021); Xu & Ziyin (2024). The main distinction between our work and prior works is that both Fukumizu (1998) and Braun et al. (2022) assumed zero-balanced ${\bf W}_{1}(0){\bf W}_{1}(0)^{T}={\bf W}_{2}(0)^{T}{\bf W}_{2}(0)$, while we relax this assumption to $\lambda$-balanced. The zero-balanced condition restricts the networks to a rich setting. We develop solutions to explore the continuum between the rich and the lazy regime. While some works, such as Tarmoun et al. (2021), have considered removing this constraint, their solutions remain in an unstable and mixed form. Our work, in its form enable the understanding of different learning regimes by exploring initialization properties beyond just absolute scale and demonstrate that this transition can be accessed and controlled by adjusting a key parameter: the relative scale. Other studies, such as Xu & Ziyin (2024) and Kunin et al. (2024), have similarly relaxed the balancedness assumption but were limited to single-output neuron settings. See Appendix A.2 for a further discussion on each of these works assumptions and their relationship to ours.
@@
As derived in Fukumizu (1998) and extended in Braun et al. (2022), whenever $\bm{F}$ is symmetric and diagonalizable such that $\bm{F}=\bm{P}\bm{\Lambda}\bm{P}^{T}$, where $\bm{P}$ is an orthonormal matrix and $\bm{\Lambda}$ is a diagonal matrix, then the unique solution to this matrix Riccatti is given by,
@@
In Appendix B.2 we prove that this equation is the unique solution to the initial value problem derived in Lemma 4.1 no matter the value of $\Lambda$. However, as discussed in Braun et al. (2022), the solution in this form is not very useable or interpretable due to the matrix inverse mixing the blocks of ${\bf Q}{\bf Q}^{T}$. Additionally, we need to diagonalize $\bm{F}$. To do so we consider the compact singular value decomposition $\operatorname{SVD}({\bf\tilde{\Sigma}}^{yx})={\bf\tilde{U}}{\bf\tilde{S}}{\bf\tilde{V}}^{T}\text{.}$ Here, $\mathbf{\tilde{U}}\in\mathbb{R}^{N_{o}\times N_{h}}$ denote the left singular vectors, $\mathbf{\tilde{S}}\in\mathbb{R}^{N_{h}\times N_{h}}$ the square matrix with ordered, non-zero eigenvalues on its diagonal, and $\mathbf{\tilde{V}}\in\mathbb{R}^{N_{i}\times N_{h}}$ the corresponding right singular vectors. For unequal input-output dimensions ($N_{i}\neq N_{o}$), the right and left singular vectors are not square. Accordingly, for the case $N_{i}>N_{h}=N_{o}$, we define $\mathbf{\tilde{U}}^{\perp}\in\mathbb{R}^{N_{o}\times|N_{o}-N_{i}|}$ as a matrix containing orthogonal column vectors that complete the basis for $\mathbf{\tilde{U}}$, i.e., make $\big{[}\mathbf{\tilde{U}}~{}\mathbf{\tilde{U}}^{\perp}\big{]}$ orthonormal, and $\mathbf{\tilde{V}}^{\perp}\in\mathbb{R}^{N_{i}\times|N_{o}-N_{i}|}$ as a matrix of zeros. Conversely, when $N_{i}=N_{h}<N_{o}$, then $\mathbf{\tilde{V}}^{\perp}$ is a matrix containing orthogonal column vectors that complete the basis for $\tilde{\bm{V}}$ and $\mathbf{\tilde{U}}^{\perp}$ is a matrix of zeros. Using this SVD structure we can now describe the eigendecomposition of $\mathbf{F}$.
@@
Under assumptions 4 and 4, the eigendecomposition of $\mathbf{F}=\mathbf{P}\bm{\Lambda}\mathbf{P}^{T}$ is
@@
Under the assumptions of whitened inputs, 4, lambda-balanced weights 4, no bottleneck 4, and full rank 4, the temporal dynamics of ${\bf Q}{\bf Q}^{T}$ are
@@
Figure 2: A The temporal dynamics of the numerical simulation (colored lines) of the loss, network function, correlation of input and output weights, and the NTK (row 1-5 respectively) are exactly matched by the analytical solution (black dotted lines) for $\lambda=-5$ B $\lambda=0.001$ Large initial weight values. C $\lambda=5$ initial weight values initialized as described in E.7.
The proof of Theorem 4.3 is in Appendix B. With this solution we can calculate the exact temporal dynamics of the loss, network function, RSMs and NTK (Fig. 2A, C) over a range of lambda-balanced initializations.
Implementation and simulation. One issue with the expression we derived in Theorem 4.3 is that it can be numerically unstable when simulating it for long time $t\gg 0$ as it involves taking the inverse of terms that involve exponentials that are diverging with $t$. If we make the additional assumption that $\bm{B}$ is invertible, then we can rearaange this expression to only use exponentials with negative coefficients, which we derive in Appendix B.5. In the next section we will discuss the significance of $\bm{B}$ being invertible at initialization on the convergence of the dynamics. Simulation details are in Appendix E.7.

### v2

# From Lazy to Rich: Exact Learning Dynamics in Deep Linear Networks

@@ Abstract
Biological and artificial neural networks develop internal representations that enable them to perform complex tasks. In artificial networks, the effectiveness of these models relies on their ability to build task-specific representations, a process influenced by interactions among datasets, architectures, initialization strategies, and optimization algorithms. Prior studies highlight that different initializations can place networks in either a lazy regime, where representations remain static, or a rich (or feature-learning) regime, where representations evolve dynamically. Here, we examine how initialization influences learning dynamics in deep linear neural networks, deriving exact solutions for $\lambda$-balanced’ initializations—defined by the relative scale of weights across layers. These solutions capture the evolution of representations and the Neural Tangent Kernel across the spectrum from the rich to the lazy regimes. Our findings deepen the theoretical understanding of the impact of weight initialization on learning regimes, with implications for continual learning, reversal learning, and transfer learning, relevant to both neuroscience and practical applications.
@@ 1 Introduction
Biological and artificial neural networks learn internal representations that enable complex tasks such as categorization, reasoning, and decision-making. Both systems often develop similar representations from comparable stimuli, suggesting shared information processing mechanisms (Yamins et al., 2014). This similarity, though not fully understood, has drawn interest from neuroscience, AI, and cognitive science (Haxby et al., 2001; Laakso & Cottrell, 2000; Morcos et al., 2018; Kornblith et al., 2019; Moschella et al., 2022). The success of neural models relies on their ability to extract relevant features from data to build internal representations, a complex process that in machine learning is defined by two regimes: lazy and rich (Saxe et al., 2014; Pennington et al., 2017; Chizat et al., 2019; Bahri et al., 2020).
@@
Lazy regime. The lazy regime follows from a fundamental phenomenon in overparameterized neural networks: during training, these networks frequently remain near their linearized form, undergoing minimal changes in the parameter space (Chizat et al., 2019). Consequently, they adopt learning dynamics akin to kernel regression, characterized by the Neural Tangent Kernel (NTK) matrix and exhibiting exponential learning behavior (Du et al., 2018; Jacot et al., 2018; Du et al., 2019; Allen-Zhu et al., 2019a; b; Zou et al., 2020). This behavior, known as the lazy or kernel regime, typically occurs in infinitely wide architectures and can be triggered by large variance initialization (Jacot et al., 2018; Chizat et al., 2019). While the lazy regime offers valuable insights into how networks converge to a global minimum, it does not fully account for the generalization capabilities of neural networks trained with standard initializations. It is, therefore, widely believed that another regime, driven by small or vanishing initializations, underpins some of the successes of neural networks.
@@
Rich regime. In contrast, the rich feature-learning regime is characterized by a NTK that evolves throughout training, accompanied by non-convex dynamics that navigate saddle points (Baldi & Hornik, 1989; Saxe et al., 2014; 2019b; Jacot et al., 2021). This regime features sigmoidal learning curves and simplicity biases, such as low-rankness (Li et al., 2020) or sparsity (Woodworth et al., 2020). Numerous studies have shown that the absolute scale of initialization drives the rich regime, which typically emerges at small initialization scales (Chizat et al., 2019; Geiger et al., 2020). However, even at small initialization scales, differences in weight magnitudes between layers can induce the lazy learning regime (Azulay et al., 2021; Kunin et al., 2024) – highlighting the significance of both absolute scale (initialization variance) and relative scale (difference in weight magnitude between layers) in generating diverse learning dynamics (Atanasov et al., 2022). Beyond absolute scale and relative scale, additional aspects of initialization can profoundly affect feature learning, including the effective rank of the weight matrices Liu et al. (2023b), layer-specific initialization variances Yang & Hu (2020); Luo et al. (2021); Yang et al. (2022), and the use of large learning rates Lewkowycz et al. (2020); Ba et al. (2022); Zhu et al. (2023); Cui et al. (2024). These findings illustrate the effect of initialization on inducing complex learning behavior through the resulting dynamics. Here we develop a solvable model which captures these diverse phenomena.
@@
Figure 1: A minimal model of the rich and lazy regimes. A. We examine a deep and wide linear network trained using gradient descent starting from an initialization characterized by a relative scale parameter $\lambda$ — which characterizes the difference in the weight covariance between the first and second layers. B. Network output for an example task over training time, starting from a range of relative scale values. The dynamics are influenced by the initialization. Solid lines represent simulations, while dotted lines indicate the analytical solutions derived in this work. C. A network with LeCun weight initialization (LeCun et al., 1998) in the infinite width limit becomes $\lambda$-balanced, as ${\bf W}_{2}^{T}{\bf W}_{2}-{\bf W}_{1}{\bf W}_{1}^{T}$ approaches the scaled identity matrix.
@@
Despite significant advances, these learning regimes and their characterization are not yet fully understood and would benefit from clearer theoretical predictions, particularly regarding the influence of prior knowledge (initialization) on the learning regime. In this work, we address this gap by deriving exact solutions for the learning dynamics in deep linear networks as a function of network initialization, providing one of the few analytical models of the rich and lazy regimes in wide and deep neural networks (Xu & Ziyin, 2024; Kunin et al., 2024; Tu et al., 2024). To illustrate the dramatic effect of initialization and the kind of phenomenon we build a theory for, we consider a two-layer linear network parameterized by an encoding layer ${\bf W}_{1}$ and a decoding layer ${\bf W}_{2}$ (Fig.1A). This network can be initialized with different relative scalings, such that ${\bf W}_{1}{\bf W}_{1}^{T}\succ{\bf W}_{2}^{T}{\bf W}_{2}$, ${\bf W}_{1}{\bf W}_{1}^{T}\ \prec{\bf W}_{2}^{T}{\bf W}_{2}$, or ${\bf W}_{1}{\bf W}_{1}^{T}={\bf W}_{2}^{T}{\bf W}_{2}$, while maintaining the same absolute scale. As shown in Fig.1B, the choice of relative scaling can result in drastically different learning trajectories and representations and the theory we develop over the course of this paper describes these effects. Through these solutions, we aim to gain insights into the rich and lazy regimes, as well as the transition between them during training, by examining the impact of relative scaling. As shown in Fig.1C and further proved in Appendix A.3, initialization methods used in practice, such as LeCun initialization in wide networks, approximate the relative scaling initialization explored in this paper, making it relevant to machine learning community as further demonstrated by Kunin et al. (2024). We consider applications relevant to machine learning and neuroscience, including continual learning (Kirkpatrick et al., 2017; Zenke et al., 2017; Parisi et al., 2019), reversal learning (Erdeniz & Atalay, 2010) and transfer learning (Taylor & Stone, 2009; Thrun & Pratt, 2012; Lampinen & Ganguli, 2018; Gerace et al., 2022).
@@
  • We derive explicit solutions for the gradient flow, internal representational similarity, and finite-width NTK in unequal-input-output two-layer deep linear networks, under a broad range of $\lambda$-balanced initialization conditions (Section 4).
@@
  • We present applications of these solutions relevant to both the neuroscience and machine learning communities, providing exact solutions for continual learning dynamics, reversal learning dynamics, and transfer learning (Section 6).
@@ 2 Related Work
Linear networks. Our work builds upon a rich body of research on deep linear networks, which, despite their simplicity, have proven to be valuable models for understanding more complex neural networks (Baldi & Hornik, 1989; Fukumizu, 1998; Saxe et al., 2014). Previous research has extensively analyzed convergence (Arora et al., 2018a; Du & Hu, 2019), generalization properties (Lampinen & Ganguli, 2018; Poggio et al., 2018; Huh, 2020), and the implicit bias of gradient descent (Arora et al., 2019a; Woodworth et al., 2020; Chizat & Bach, 2020; Kunin et al., 2022) in linear networks. These studies have also revealed that deep linear networks have intricate fixed point structures and nonlinear learning dynamics in parameter and function space, reminiscent of phenomena observed in nonlinear networks (Arora et al., 2018b; Lampinen & Ganguli, 2018). Seminal work by Saxe et al. (2014) laid the groundwork by providing exact solutions to gradient flow dynamics under task-aligned initializations, demonstrating that the largest singular values are learned first during training. This analysis has been extended to deep linear networks (Arora et al., 2018b; 2019a; Ziyin et al., 2022) with more flexible initialization schemes (Gidel et al., 2019; Tarmoun et al., 2021; Gissin et al., 2019). This work directly builds on the matrix Riccati formulation proposed by Fukumizu (1998) and Braun et al. (2022) which extends these solutions to wide networks. We extend and refine these results to obtain the dynamics for a wider class of $\lambda$-balanced networks to more clearly demonstrate the impact of initialization on rich and lazy learning regimes also developed in Tu et al. (2024) for a set of orthogonal initalizations. Our work extends previous analyses (Xu & Ziyin, 2024; Kunin et al., 2024) of these regimes to wide networks. Previous studies leveraged these solutions primarily to characterize convergence rates; however, our work goes beyond this by providing a comprehensive characterization of the complete dynamics of the system (Tarmoun et al., 2021).
@@
Infinite-width networks. Recent advances in understanding the rich regime have largely stemmed from examining how the initialization variance and layer-wise learning rates must scale in the infinite-width limit to maintain consistent behavior in activations, gradients, and outputs. Several studies have employed statistical mechanics tools to derive analytical solutions for the rich population dynamics of two-layer nonlinear neural networks initialized using a mean-field parameterization (Mei et al., 2018; Rotskoff & Vanden-Eijnden, 2018; Chizat & Bach, 2018; Sirignano & Spiliopoulos, 2020; Rotskoff & Vanden-Eijnden, 2022; Sirignano & Spiliopoulos, 2020). Other methods for analyzing deep network dynamics include the NTK limit, where the network effectively performs kernel regression without feature learning (Jacot et al., 2018; Lee et al., 2019; Arora et al., 2019b). Furthermore, these approaches typically require numerical integration or operate within a limited learning regime, and are unable to describe the learning dynamics of hidden representations. Instead, our work provides explicit analytical solutions for the dynamics of the network and its NTK in the finite-width case (Jacot et al., 2021; Xu & Ziyin, 2024; Kunin et al., 2024; Chizat et al., 2019).
@@ 3 Preliminaries
and the initialization ${\bf W}_{2}(0)$,${\bf W}_{1}(0)$. We employ an approach first introduced in the foundational work of Fukumizu (1998) and extended in recent work by Braun et al. (2022), which instead of studying the parameters directly, considers the dynamics of a matrix of the important statistics. In particular, defining ${\bf Q}=\begin{bmatrix}{\bf W}_{1}&{\bf W}_{2}^{T}\end{bmatrix}^{T}\in\mathbb{R}^{(N_{i}+N_{o})\times N_{h}}$, we consider the $(N_{i}+N_{o})\times(N_{i}+N_{o})$ matrix
@@
where $+$ denotes the pseudoinverse; and the network’s finite-width NTK (Jacot et al., 2018; Lee et al., 2019; Arora et al., 2019b)
@@
where ${\bf I}_{N_{o}}$ is the $N_{o}\times N_{o}$ identity matrix and $\otimes$ is the Kronecker product. Hence, the dynamics of ${\bf Q}{\bf Q}^{T}$ describes the important aspects of network behaviour.
@@ 4 Exact Learning Dynamics
We derive an exact solution for ${\bf Q}{\bf Q}^{T}$ offering insight into the learning dynamics, convergence behavior, and generalization properties of two-layer linear networks with prior knowledge.
@@
  • A1 (Whitened input). The input data is whitened, i.e. ${\bf\tilde{\Sigma}}^{xx}={\bf I}$.
  • A2 ($\lambda$-Balanced). The network’s weight matrices are $\lambda$-balanced at the beginning of training, i.e. ${\bf W}_{2}^{T}{\bf W}_{2}(0)-{\bf W}_{1}{\bf W}_{1}(0)^{T}=\lambda\mathbf{I}$. If this condition holds at initialization, it will persist throughout training (Saxe et al., 2014; Arora et al., 2018a). For completeness, we prove this in Appendix A.
@@
These assumptions are strictly weaker than prior works (Fukumizu, 1998; Braun et al., 2022; Kunin et al., 2024; Xu & Ziyin, 2024). The main distinction between our work and prior works is that both Fukumizu (1998) and Braun et al. (2022) assumed zero-balanced weights (${\bf W}_{1}(0){\bf W}_{1}(0)^{T}={\bf W}_{2}(0)^{T}{\bf W}_{2}(0)$), while we relax this assumption to $\lambda$-balanced. The zero-balanced condition restricts the networks to a rich setting. We develop solutions to explore the continuum between the rich and the lazy regime. While some works, such as Tarmoun et al. (2021), have considered removing this constraint, their solutions remain in an unstable and mixed form. Other studies, such as Xu & Ziyin (2024) and Kunin et al. (2024), have similarly relaxed the balanced assumption but were limited to single-output neuron settings. See Appendix A.2 for a further discussion on each of these works’ assumptions and their relationship to ours.
@@
As derived in Fukumizu (1998) and extended in Braun et al. (2022), whenever $\bm{F}$ is symmetric and diagonalizable such that $\bm{F}=\bm{P}\bm{\Lambda}\bm{P}^{T}$, where $\bm{P}$ is an orthonormal matrix and $\bm{\Lambda}$ is a diagonal matrix, then the unique solution to this matrix Riccatti equation is given by
@@
In Appendix B.2 we prove that this equation is the unique solution to the initial value problem derived in Lemma 4.1 for any value of $\Lambda$. However, as discussed in Braun et al. (2022), the solution in this form is not very useable or interpretable due to the matrix inverse mixing the blocks of ${\bf Q}{\bf Q}^{T}$. Additionally, we need to diagonalize $\bm{F}$. To do so we consider the compact singular value decomposition $\operatorname{SVD}({\bf\tilde{\Sigma}}^{yx})={\bf\tilde{U}}{\bf\tilde{S}}{\bf\tilde{V}}^{T}\text{.}$ Here, $\mathbf{\tilde{U}}\in\mathbb{R}^{N_{o}\times N_{h}}$ denotes the left singular vectors, $\mathbf{\tilde{S}}\in\mathbb{R}^{N_{h}\times N_{h}}$ the square matrix with ordered, non-zero eigenvalues on its diagonal, and $\mathbf{\tilde{V}}\in\mathbb{R}^{N_{i}\times N_{h}}$ the corresponding right singular vectors. For unequal input-output dimensions ($N_{i}\neq N_{o}$), the right and left singular vectors are not square. Accordingly, for the case $N_{i}>N_{h}=N_{o}$, we define $\mathbf{\tilde{U}}_{\perp}\in\mathbb{R}^{N_{o}\times|N_{o}-N_{i}|}$ as a matrix containing orthogonal column vectors that complete the basis for $\mathbf{\tilde{U}}$, i.e., make $\big{[}\mathbf{\tilde{U}}~{}\mathbf{\tilde{U}}_{\perp}\big{]}$ orthonormal, and $\mathbf{\tilde{V}}_{\perp}\in\mathbb{R}^{N_{i}\times|N_{o}-N_{i}|}$ as a matrix of zeros. Conversely, when $N_{i}=N_{h}<N_{o}$, then $\mathbf{\tilde{V}}_{\perp}$ is a matrix containing orthogonal column vectors that complete the basis for $\tilde{\bm{V}}$ and $\mathbf{\tilde{U}}_{\perp}$ is a matrix of zeros. Using this SVD structure we can now describe the eigendecomposition of $\mathbf{F}$.
@@
Under assumptions 4, the eigendecomposition of $\mathbf{F}=\mathbf{P}\bm{\Lambda}\mathbf{P}^{T}$ is
@@
Under the assumptions of whitened inputs (4), $\lambda$-balanced weights (4), and no bottleneck (4), the temporal dynamics of ${\bf Q}{\bf Q}^{T}$ are
@@
Figure 2: A. The temporal dynamics of the numerical simulation (colored lines) of the loss, network function, correlation of input and output weights, and the NTK (row 1-5 respectively) are exactly matched by the analytical solution (black dotted lines) for $\lambda=-2$. B. $\lambda=0$ Large initial weight values. C. $\lambda=2$ initial weight values initialized as described in E.7.
The proof of Theorem 4.3 is in Appendix B. With this solution we can calculate the exact temporal dynamics of the loss, network function, RSMs and NTK (Fig. 2A, C) over a range of $\lambda$-balanced initializations.
Implementation and simulation. One issue with the expression we derived in Theorem 4.3 is that it can be numerically unstable when simulating it for long time $t\gg 0$ as it involves taking the inverse of terms that involve exponentials that are diverging with $t$. If we make the additional assumption that $\bm{B}$ is invertible, then we can rearrange this expression to only use exponentials with negative coefficients, which we derive in Appendix B.5. In the next section we will discuss the significance of $\bm{B}$ being invertible at initialization on the convergence of the dynamics. Simulation details are in Appendix E.7.